### v1

# Text-Guided Layer Fusion Mitigates Hallucination in Multimodal LLMs

Abstract
Multimodal large language models (MLLMs) typically rely on a single late-layer feature from a frozen vision encoder, leaving the encoder’s rich hierarchy of visual cues under-utilized. MLLMs still suffer from visually ungrounded hallucinations, often relying on language priors rather than image evidence. While many prior mitigation strategies operate on the text side, they leave the visual representation unchanged and do not exploit the rich hierarchy of features encoded across vision layers. Existing multi-layer fusion methods partially address this limitation but remain static, applying the same layer mixture regardless of the query. In this work, we introduce TGIF (Text-Guided Inter-layer Fusion), a lightweight module that treats encoder layers as depth-wise “experts” and predicts a prompt-dependent fusion of visual features. TGIF follows the principle of direct external fusion, requires no vision-encoder updates, and adds minimal overhead. Integrated into LLaVA-1.5-7B, TGIF provides consistent improvements across hallucination, OCR, and VQA benchmarks, while preserving or improving performance on ScienceQA, GQA, and MMBench. These results suggest that query-conditioned, hierarchy-aware fusion is an effective way to strengthen visual grounding and reduce hallucination in modern MLLMs. Our code will be available at: https://github.com/Linchenchen/TGIF. Refer to caption Figure 1: Comparison of layer fusion designs in MLLMs. (a) MLP Connector: Uses only the penultimate layer of the vision encoder, mapping global visual tokens through a simple projection. (b) Dense Connector: Aggregates multi-layer visual features via concatenation or downsampling before projection, enriching semantics but with a fixed fusion pattern. (c) MMFuser: Retrieves shallow-layer features using deep-layer queries (Q–K/V attention) to capture local details. (d) Proposed TGIF: Introduces a text-guided layer router that dynamically reweights features from multiple layers based on the input query, enabling adaptive, context-aware visual fusion.
1 Introduction
Multimodal large language models (MLLMs) have recently achieved impressive progress on visual question answering, captioning, and open-ended multimodal dialogue by combining the reasoning ability of large language models (LLMs) with the perceptual capacity of pretrained vision encoders [llava, li2023blip, instructblip]. Most state-of-the-art systems adopt a modular design: a frozen vision encoder (e.g., CLIP ViT), a lightweight connector, and a powerful LLM decoder. The connector projects visual embeddings into the LLM’s token space, enabling the model to jointly process image and text tokens for downstream reasoning.
Despite these advances, modern MLLMs still frequently produce confident but visually ungrounded descriptions, a phenomenon broadly referred to as hallucination [rohrbach2019objecthallucinationimagecaptioning, liu2024surveyhallucinationlargevisionlanguage]. In the vision–language setting, hallucination typically manifests as objects, attributes, or relations that are plausible under language priors but inconsistent with the input image. This issue is especially severe for detail-oriented tasks (e.g., OCR, small object recognition), where high-level semantic features alone are insufficient to support fine-grained grounding.
Prior work tackles hallucination from two main angles. Training-based approaches improve alignment via additional instruction tuning, contrastive fine-tuning, or RLHF [gunjal2024detectingpreventinghallucinationslarge, yu2024rlhfvtrustworthymllmsbehavior], but they are often data- and compute-intensive. Training-free approaches instead modify decoding or inference without retraining the model. Methods such as VCD [leng2023mitigatingobjecthallucinationslarge], VTI [liu2024reducinghallucinationsvisionlanguagemodels], OPERA [huang2024opera], FarSight [tang2025seeingfarclearlymitigating], and PerturboLLaVA [chen2025perturbollavareducingmultimodalhallucinations] mitigate hallucination via contrastive calibration, latent intervention, causal masking, or over-trust penalties. However, these strategies primarily operate on the text side: they adjust the decoder or filter responses post hoc, while the underlying visual representation typically remains a single, fixed layer of the vision encoder passed through an MLP projector.
Meanwhile, transformer-based vision encoders such as CLIP are known to build a rich hierarchy of visual abstractions across layers: shallow layers preserve edges, textures, and local geometry, whereas deeper layers encode high-level semantics aligned with text [clip, gandelsman2024interpretingclipsimagerepresentation]. Recent works have begun to exploit this hierarchy by fusing features from multiple depths. DenseConnector [yao2024denseconnectormllms] concatenates or downsamples features from selected layers, and MMFuser [cao2024mmfusermultimodalmultilayerfeature] retrieves shallow-layer details using deep features as queries. A recent systematic study further shows that direct, external fusion of multi-layer visual features at the input stage yields the most stable performance across architectures [lin2025multilayervisualfeaturefusion]. However, existing fusion schemes are static: the mixture of layers is fixed once chosen, independent of the question, and tends to overemphasize globally aligned semantics even when the query requires local evidence.
In this work, we argue that hallucination in MLLMs is fundamentally linked to how visual features are selected and exposed to the LLM. We propose TGIF (Text-Guided Inter-layer Fusion), a dynamic routing framework that treats the layers of a frozen vision encoder as a pool of specialized “experts” and adaptively fuses them based on the input query. Concretely, TGIF introduces a lightweight router inside the multimodal projector that takes text (and optionally a global image feature) as input and outputs a soft distribution over all ViT layers. These learned weights are then used to form a query-conditioned fused visual representation, which is projected into the LLM token space using a standard MLP connector (Fig. 2). Following the principle of direct external fusion [lin2025multilayervisualfeaturefusion], TGIF keeps the vision encoder frozen and preserves the token budget, while allowing the layer mixture to vary per prompt.
We instantiate TGIF on top of LLaVA-1.5 [li2023llava] with CLIP-ViT-L/14 and Vicuna-7B, and evaluate it across three benchmark families: hallucination, OCR, and general VQA. On hallucination-focused evaluations, TGIF achieves state-of-the-art performance among 7B-scale LLaVA variants, improving POPE accuracy from 86.85% to 87.91% and boosting HallusionBench All Accuracy from 46.90% to 49.94%, outperforming both decoding-based baselines (VCD, OPERA, VTI, FarSight, PerturboLLaVA) and larger 13B models. On OCRBench, TGIF improves the final score from 297 to 313 (+16), and provides consistent gains on TextVQA. At the same time, it maintains or improves overall performance on ScienceQA, GQA, and MMBench, indicating that better grounding does not come at the expense of high-level reasoning.
Our main contributions are three-fold:
  • We identify a key limitation of current multimodal LLMs: visual tokens are typically drawn from a single, late-layer representation, which is poorly suited for detail-sensitive grounding and exacerbates hallucination under strong language priors.
  • We propose TGIF, a text-guided inter-layer fusion module that dynamically reweights CLIP layers per query, following the direct external fusion principle while remaining parameter- and token-efficient. We explore both text-only and multimodal routers, and introduce an entropy-based load-balancing loss to prevent expert collapse.
  • We demonstrate that TGIF substantially improves hallucination robustness and fine-grained visual perception on POPE, HallusionBench, OCRBench, and TextVQA, while preserving competitive performance on general reasoning benchmarks. Qualitative analysis of router behavior further shows that TGIF learns semantically meaningful depth-selection patterns across task types.
2 Related Work
2.1 Multimodal Large Language Models
Multimodal large language models (MLLMs) integrate the reasoning ability of LLMs with the perceptual capacity of pretrained vision encoders. Most follow a modular design with a frozen vision encoder, a lightweight connector, and a powerful LLM decoder. The connector projects visual embeddings into the text space, enabling cross-modal alignment essential for grounded reasoning. Early designs used simple MLP projectors [llava], while later approaches like BLIP-2 [li2023blip] and InstructBLIP [instructblip] employ query-based modules for salient visual token extraction.
2.2 Hallucination in MLLMs
MLLMs frequently generate confident yet ungrounded content—a phenomenon known as hallucination [rohrbach2019objecthallucinationimagecaptioning, liu2024surveyhallucinationlargevisionlanguage]. Training-based mitigation strategies rely on additional instruction tuning, contrastive fine-tuning, or RLHF [gunjal2024detectingpreventinghallucinationslarge, yu2024rlhfvtrustworthymllmsbehavior], but they are compute- and data-intensive. Training-free methods instead modify decoding or inference. Recent works such as VCD [leng2023mitigatingobjecthallucinationslarge], VTI [liu2024reducinghallucinationsvisionlanguagemodels], OPERA [huang2024opera], FarSight [tang2025seeingfarclearlymitigating], and PerturboLLaVA [chen2025perturbollavareducingmultimodalhallucinations] mitigate hallucination via contrastive calibration, token intervention, or causal masking. However, these approaches primarily act on the text side, leaving the underlying vision-language alignment unchanged. Our work complements them by targeting hallucination at the feature level through text-guided multi-layer fusion.
2.3 Multi-Layer Visual Feature Fusion
Transformer-based vision encoders like CLIP exhibit a hierarchical structure where deeper layers capture semantic abstraction and intermediate layers preserve fine-grained spatial cues [clip, gandelsman2024interpretingclipsimagerepresentation]. DenseConnector [yao2024denseconnectormllms] concatenates features from multiple layers, while MMFuser [cao2024mmfusermultimodalmultilayerfeature] retrieves shallow-layer details using deep-layer queries. These static strategies enrich representations but cannot adapt fusion to each query. Recent work also systematically analyzes layer integration strategies, showing that direct external fusion yields the most stable performance [lin2025multilayervisualfeaturefusion]. Building on these insights, TGIF performs text-guided inter-layer fusion, dynamically reweighting visual features according to the input query to improve grounding and reduce hallucination.
3 Method
Our approach is situated within a standard Vision Language Model (VLM) architecture, which comprises a vision encoder (e.g., CLIP ViT [clip]), a language model (e.g., Vicuna [vicuna]), and a multimodal projector that maps visual features into the language model’s embedding space. Our proposed text-guided layer selection module is designed as a core component of this multimodal projector. It takes as input the full stack of hidden states from the all vision encoder layers, along with features derived from the text prompt, to produce a dynamically tailored visual representation for the LLM.
Figure 2: Overview of the proposed Text-Guided Inter-Layer Fusion (TGIF) framework. TGIF dynamically integrates hierarchical visual features from a frozen vision encoder based on the textual query. The image is first processed by the Vision Transformer (ViT), producing multi-layer representations $\{F_{l}\}$ that capture progressively abstract semantics. The Layer Router receives the text embedding $f_{\text{text}}$ and outputs a soft distribution over encoder layers $w_{\text{layer}}$ through an MLP and softmax. These weights determine the contribution of each layer to the fused visual feature $F_{\text{fused}}$, which is then projected to the text space by a lightweight MLP connector. The fused multimodal tokens are concatenated with the tokenized text and fed into the LLM for reasoning and response generation.
3.1 Text-Guided Layer Selection
As established in prior work [chen2025multimodallanguagemodelsbetter], different CLIP layers capture distinct semantic information, with shallow layers encoding textures and spatial details while deeper layers align more with global semantics. This suggests that the layers can be viewed as a pool of specialized “experts”. Borrowing concepts from the Mixture-of-Experts (MoE) paradigm, we treat each layer as an expert and propose a dynamic layer selection framework, which we name TGIF (Text-Guided Inter-layer Fusion). In this framework, a text-guided “router” learns to generate weights to select and fuse the most relevant layer experts for a given task. We explore two architectures for the TGIF router, one with text-only input and other with multimodal input. The model architecture and framework is shown in Fig 2. This design allows adaptive, query-conditioned fusion of depth-wise visual cues, improving both grounding and fine-grained detail understanding.
3.1.1 Text-Guided MLP Router
Our baseline approach uses a lightweight MLP-based router to predict layer importance scores based solely on the textual prompt. This router learns a direct mapping from the question’s semantics to the relevance of different vision layers.
Let $\mathbf{f}_{\text{text}}\in\mathbb{R}^{D_{t}}$ denote the pooled text embedding from the LLM. Let $\{\mathbf{F}_{l}\in\mathbb{R}^{P\times D_{v}}\}_{l=1}^{L}$ be the set of patch-level visual features from all $L$ layers of the vision encoder. The MLP selector first predicts unnormalized logits for each layer:
$\mathbf{z}=\text{MLP}(\mathbf{f}_{\text{text}})\in\mathbb{R}^{L}.$ (1)
These logits are then transformed into a probability distribution using the softmax function to represent the layer weights:
$\boldsymbol{w}=\text{softmax}(\mathbf{z})\in\mathbb{R}^{L}.$ (2)
The final fused visual representation is computed as a weighted sum of all layer features, where the weights are broadcast across the patch and feature dimensions:
$\mathbf{F}_{\text{fused}}=\sum_{l=1}^{L}w_{l}\cdot\mathbf{F}_{l}\in\mathbb{R}^{P\times D_{v}}.$ (3)
3.1.2 Multimodal MLP Router
To address cases in pretraining where the text prompt is generic (e.g., “Describe the image”), we enhance the router with visual context. This multimodal approach allows the selection to be conditioned on both the question and the image content.
We first extract a global image representation, $\mathbf{f}_{\text{image}}\in\mathbb{R}^{D_{v}}$, by taking the [CLS] token from the penultimate layer of the vision encoder. The text and image features are then projected to a common dimension $D_{p}$ and concatenated:
$\mathbf{f}_{\text{multi}}=[\mathbf{f}_{\text{text}}\mathbf{W}_{t},\mathbf{f}_{\text{image}}\mathbf{W}_{v}]\in\mathbb{R}^{2D_{p}},$ (4)
where $\mathbf{W}_{t}\in\mathbb{R}^{D_{t}\times D_{p}}$ and $\mathbf{W}_{v}\in\mathbb{R}^{D_{v}\times D_{p}}$ are learnable projection matrices. This combined multimodal feature vector is then used by the MLP to predict the layer weights, following the same process as in Equations 1-3.
3.2 Load Balancing Loss
A common challenge when training MoE-style routers is the tendency for the router to converge to a state where it consistently selects the same few “safe” experts (in our case, layers), leading to “expert starvation” [shazeer2017outrageouslylargeneuralnetworks]. To mitigate this and encourage the router to utilize a more diverse set of layers, we incorporate a modified auxiliary load balancing loss into our total training objective.
For our soft-selection routers, we use an entropy-based loss. Let $\boldsymbol{w}_{b}\in\mathbb{R}^{L}$ be the layer weights for the $b$-th sample in a batch of size $B$. We first compute the average weight for each layer across the batch:
$\bar{\boldsymbol{w}}=\frac{1}{B}\sum_{b=1}^{B}\boldsymbol{w}_{b}\in\mathbb{R}^{L}.$ (5)
The auxiliary loss is then formulated to maximize the entropy of this average distribution, which encourages a more uniform usage of layers:
$\mathcal{L}_{\text{aux}}=\lambda\sum_{l=1}^{L}\bar{w}_{l}\log(\bar{w}_{l}+\epsilon),$ (6)
where $\lambda$ is a hyperparameter controlling the strength of the loss and $\epsilon$ is a small constant for numerical stability. This auxiliary loss is added to the main VLM loss during training.
Because the nature of textual prompts differs between pretraining and instruction tuning dataset, we apply different $\lambda$ values across stages. During pretraining, the router often receives generic prompts (e.g., “Describe the image”), which provide limited textual guidance. We thus strengthen the visual signal by applying a slightly larger $\lambda$ to encourage exploration of multiple layers. During fine-tuning, prompts are task-oriented and semantically rich (e.g., “What number is written on the sign?”). Here, we reduce $\lambda$ to allow the router to focus on discriminative, text-conditioned layer selection.
4 Experiments
Method Hallucination OCR Overall
POPE HallusionBench TextVQA OCRBench ScienceQA GQA MMBench
Baseline Models
LLaVA-1.5-7B 86.85 46.27 58.20 30.80 66.80 62.00 64.30
+ Dense Connector 86.60 59.20 69.50 63.80 66.80
+ MMFuser 86.30 58.80 68.70 62.80 67.50
Proposed Methods (TGIF)
+ TGIF (Text-Only MLP) 87.30 49.95 58.93 31.50 68.07 62.48 65.97
+ TGIF (Multimodal MLP) 86.26 57.31 59.09 29.90 69.72 62.37 65.46
+ TGIF ($\lambda{=}0.01$ pretrain-only) 87.91 48.68 58.98 31.30 70.10 62.58 66.40
Table 1: Comparison across VLM benchmarks. Hallucination Benchmarks: POPE and HallusionBench evaluate factual grounding and object hallucination. OCR Benchmarks: TextVQA and OCRBench assess text recognition and fine-grained perception. Overall Performance: ScienceQA, GQA, and MMBench measure general reasoning and instruction-following. Best per metric in bold; second best underlined.
Method Accuracy $\uparrow$ F1 Score $\uparrow$
LLaVA-1.5-7B 86.85 85.86
+ VCD 84.66 84.51
+ OPERA 84.20 85.40
+ VTI 86.50 85.90
+ FarSight 86.10 80.40
+ TGIF (ours) 87.91 86.23
Table 2: POPE Comparison. We report the average F1-score and accuracy averaged across three sub-tasks. Detailed results are provided in the Appendix.
Method Params All Acc. $\uparrow$
LLaVA-1.5 13.0B 46.94
Qwen-VL 9.6B 39.15
Open-Flamingo 9.0B 38.44
InstructBLIP 8.2B 45.26
MiniGPT5 8.2B 40.30
MiniGPT4 8.2B 35.78
LLaVA-1.5 7.0B 46.90
+ VCD 7.0B 46.90
+ OPERA 7.0B 47.10
+ PerturboLLaVA 7.0B 47.60
+ TGIF (ours) 7.0B 49.94
Table 3: HallusionBench Comparison. We compare LLaVA-1.5-7B+TGIF with similarly sized open-source models (7–13B). We report GPT4-assisted All Accuracy. Detailed results are provided in the Appendix.
In this section, we present a comprehensive evaluation of our proposed TGIF framework. We compare our best-performing model against the LLaVA-1.5 baseline and then provide a detailed analysis of our ablation studies to understand the contributions of different components of our design.
4.1 Experimental Setting
4.1.1 Implementation Details
We implement our proposed framework, TGIF, on top of the publicly available LLaVA-1.5 [li2023llava] codebase. To ensure a fair comparison, our primary experiments maintain consistency with LLaVA-1.5 by employing CLIP-ViT-L/14-336px [clip] as the vision encoder and Vicuna-7B [vicuna] as the LLM.
4.1.2 Training Recipe
We train all models on 8 NVIDIA H100-80G GPUs following the two-stage training paradigm of LLaVA-1.5 [li2023llava].
Stage 1 (Feature Alignment Pretraining). In this stage, both the vision encoder and the LLM remain frozen. We train only the TGIF components, the layer router and MLE connector module, on a filtered subset of 558K image–text pairs from CC3M. This stage aligns multi-layer visual representations with the LLM’s embedding space and allows the router to learn an initial prompt-adaptive layer selection policy. We use a learning rate of 1e-3 and a global batch size of 256.
Stage 2 (Instruction Finetuning). Following pretraining, we fine-tune the LLM together with the TGIF projector while keeping the vision encoder frozen. Here the dataset we use is the 665K multi-turn conversations for instruction fine-tuning. This phase refines the model’s conversational and reasoning abilities using dynamically fused visual features. The learning rate is set to 2e-5 with a batch size of 128.
We also apply stage-specific load balancing coefficients (Sec. 3.2) to encourage diverse layer usage during pretraining and more discriminative text-conditioned routing during fine-tuning.
4.1.3 Evaluation Benchmarks
To comprehensively evaluate our TGIF framework, we benchmark its performance across a diverse set of multimodal tasks spanning hallucination detection, fine-grained OCR reasoning, and general visual question answering. All evaluations are conducted using the standardized VLMEvalKit [duan2024vlmevalkit] platform to ensure consistency and comparability across models.
Hallucination Benchmarks. To assess grounding faithfulness, we adopt two representative hallucination benchmarks: HallusionBench (HB) [hallusionbench] and POPE [Li-hallucination-2023]. HallusionBench probes visual factuality by testing a model’s ability to reject implausible object claims, while POPE reformulates hallucination detection as a binary classification task, quantifying the model’s awareness of object existence.
OCR Benchmarks. To evaluate text recognition and detail-sensitive reasoning, we include TextVQA [singh2019vqamodelsread] and the comprehensive OCRBench [Liu_2024], which covers scene-text, document-text, and key information extraction subtasks. These benchmarks reveal the model’s ability to retrieve and interpret fine-grained textual cues—an area where hallucination often manifests due to overreliance on semantic priors.
General Reasoning and Overall Benchmarks. For overall multimodal reasoning performance, we report results on widely used visual QA and instruction-following datasets, including MMBench (MMB) [liu2024mmbench], ScienceQA [lu2022learn] and GQA [ainslie2023gqatraininggeneralizedmultiquery]. Together, these benchmarks measure both the factual grounding and generalization capability of TGIF across visual domains and task types.
4.1.4 Baseline Methods
We evaluate TGIF against two complementary categories of baselines: (1) hallucination-mitigation methods that focus on decoding and inference, and (2) multi-layer fusion architectures that enhance visual representations.
Hallucination Mitigation Methods. We compare TGIF with five representative training-free approaches that operate on the decoding side. VCD [leng2023mitigatingobjecthallucinationslarge] introduces visual contrastive decoding, contrasting outputs from original and distorted inputs to reduce unimodal bias. VTI [liu2024reducinghallucinationsvisionlanguagemodels] stabilizes cross-modal interactions by injecting visual and textual interventions in latent space during inference. OPERA [huang2024opera] applies an over-trust penalty and retrospection-allocation decoding to discourage over-attention to summary tokens. FarSight [tang2025seeingfarclearlymitigating] improves token propagation through causal masking, mitigating attention drift toward outlier tokens. PerturboLLaVA [chen2025perturbollavareducingmultimodalhallucinations] enhances robustness by perturbing the visual embedding space during training to counteract language priors. These methods primarily target textual hallucinations during generation and provide a strong benchmark for comparison on HallusionBench and POPE.
Layer Fusion Baselines. For fair evaluation of visual representation improvements, we also compare with Dense Connector [yao2024denseconnectormllms] and MMFuser [cao2024mmfusermultimodalmultilayerfeature], which aggregate visual information from multiple layers of the vision encoder. Dense Connector concatenates or downsamples features from selected depths, while MMFuser retrieves shallow-layer details using deep features as queries. These architectures serve as strong baselines for evaluating how TGIF’s dynamic, text-guided fusion enhances multimodal grounding and general reasoning.
Method Recog. $VQA^{S}$ $VQA^{D}$ KIE HMER Final
LLaVA1.5-7B 160 117 15 5 0 297
TGIF (ours) 162 121 24 6 0 313
Table 4: OCRBench Comparison. Recog.:Text Recognition; $VQA^{S}$:Scene Text-centric VQA; $VQA^{D}$:Document-oriented VQA; KIE: Key Information Extraction; HMER:Handwritten Math Expression Recognition. TGIF improves over LLaVA-1.5-7B by +16 on the Final score.
4.2 Experimental Results
Table 1 summarizes TGIF’s quantitative performance across three evaluation families: hallucination, OCR, and general reasoning. Our method consistently improves fine-grained grounding and text perception while maintaining strong overall reasoning ability. Compared to static layer-fusion baselines such as DenseConnector and MMFuser, TGIF’s text-guided routing achieves clear gains on hallucination and OCR benchmarks, validating the effectiveness of adaptive, query-aware fusion. TGIF delivers consistent improvement on hallucination-focused benchmarks (+3.7% on HallusionBench, +1.1% on POPE) and OCR tasks (+0.9% on TextVQA, +0.7% on OCRBench), while matching or surpassing DenseConnector and MMFuser on overall reasoning benchmarks. These results demonstrate that text-guided inter-layer fusion enhances grounding precision without compromising high-level semantics.
We further analyze TGIF’s performance on hallucination mitigation, fine-grained perception, and general reasoning to understand its behavior across tasks. For detailed performance comparisons, we adopt the best-performing configuration, TGIF with a pretraining-only load-balancing coefficient of $\lambda=0.01$.
4.2.1 Hallucination mitigation
On POPE (Table 2), TGIF achieves the highest accuracy (87.91%) and F1 score (86.23%), surpassing recent decoding-based methods including VCD, OPERA, VTI, and FarSight. On HallusionBench (Table 3), TGIF attains an All Accuracy of 49.94%, outperforming LLaVA-1.5 by +3.0% and exceeding larger 13B-parameter models. This indicates that TGIF reduces both structured and generative hallucinations by providing richer, query-conditioned visual grounding.
4.2.2 Fine-grained perception
TGIF strengthens visual-text alignment on text-centric reasoning tasks. As shown in Table 4, TGIF improves the overall OCRBench score by +16 points over LLaVA-1.5-7B, driven by better recognition and document VQA accuracy. The gains mainly stem from TGIF’s ability to emphasize low- to mid-level layers that encode edges, text strokes, and local layout cues—features often overlooked by single-layer connectors. These improvements confirm TGIF’s suitability for dense and detail-sensitive multimodal tasks.
4.2.3 General reasoning
Across general benchmarks (ScienceQA, GQA, MMBench), TGIF maintains competitive reasoning ability. The $\lambda{=}0.01$ pretrain-only variant achieves the best ScienceQA accuracy (70.1%) and a strong 66.4% MMBench score, suggesting that dynamic layer fusion generalizes well to unseen instructions. Slight variation in GQA performance is attributed to TGIF’s stronger grounding, which prioritizes factual consistency over speculative generation. Overall, TGIF acts as a grounding-aware regularizer, improving trustworthiness without sacrificing general reasoning.
4.3 Discussion
4.3.1 Router Layer Selection Dynamics
Figure 3: Router layer selection patterns across different question categories. This heatmap visualizes the router’s learned weights for selecting vision transformer (ViT) layers across three categories of questions: General, Hallucination Detection, and OCR/Detail Recognition. Each row corresponds to one question, and each column indicates a specific ViT layer. Brighter colors denote higher selection weight for that layer.
To understand how TGIF adapts visual fusion to different query types, we visualize the learned layer-weight distributions in Fig. 3. The router exhibits clear, semantically-driven routing patterns. General queries (e.g., “Describe the image.”) activate a broad mixture of mid- and high-level layers, reflecting the need for holistic scene understanding. Hallucination-sensitive queries place greater weight on early layers that preserve spatial and boundary cues, which helps verify object presence rather than relying on language priors. In contrast, OCR and detail-oriented questions concentrate weight on mid-to-late layers containing rich text strokes and structural detail. These behaviors confirm that TGIF does not rely on a fixed mixture but instead performs question-aware selection of visual experts, addressing the limitations of static multi-layer fusion.
4.3.2 Impact of Multimodal Guidance
We compare the Text-Only and Multimodal versions of our router. Incorporating a global visual token consistently improves performance on benchmarks requiring precise grounding, such as HallusionBench, POPE, and TextVQA. Visual context helps the router disambiguate whether the query demands global semantics or local fine-grained evidence, enabling more accurate layer reweighting. This validates our hypothesis that multimodal guidance strengthens routing decisions beyond what text alone can provide.
4.3.3 Effect of Load-balancing Loss
[…]
Figure 4: Effect of load balancing on the VQA–hallucination trade-off. Each point shows the average VQA score (ScienceQA, GQA, TextVQA) versus POPE accuracy for a different router / load-balancing configuration. We annotate all load-balancing settings next to their corresponding points.
We further analyze the impact of the entropy-based load-balancing loss on routing stability and downstream performance. As shown in Fig. 4, applying a small amount of regularization during pretraining only provides the best balance between VQA accuracy and hallucination robustness. In particular, the $\lambda{=}0.01$ pretrain-only configuration yields both the highest POPE accuracy and the strongest average VQA score, indicating that mild early-stage entropy encourages the router to explore a diverse set of layers without suppressing its ability to specialize.
In contrast, larger coefficients (e.g., $\lambda{=}0.1$) or applying the loss throughout full fine-tuning tend to over-regularize the router, pulling the layer distribution toward uniformity and reducing its ability to adapt the fusion pattern to the input query. These settings achieve weaker VQA performance and, in some cases, degraded hallucination resistance. Overall, the results suggest that light, pretraining-only regularization is crucial: it stabilizes routing and prevents expert collapse while still allowing the model to learn query-dependent, discriminative layer selection during instruction tuning.
5 Conclusion
In this work, we revisited a core assumption in multimodal large language models: the use of a single deep-layer visual representation as the primary input to the LLM. Our study shows that this design restricts fine-grained grounding and increases the likelihood of hallucination. To address this limitation, we introduced TGIF, a lightweight and training-efficient module that performs text-guided inter-layer fusion over a frozen vision encoder. TGIF treats encoder layers as depth-wise experts and routes them according to the input query, allowing the LLM to access richer, hierarchy-aware visual information without modifying the token budget or updating the vision encoder.
Comprehensive experiments across hallucination benchmarks, OCR tasks, and general VQA confirm that TGIF improves grounding quality and reduces hallucination. The method achieves state-of-the-art results among 7B-scale LLaVA variants and even surpasses larger models on HallusionBench. An analysis of router behaviors shows that TGIF produces semantically meaningful depth-selection patterns: it prioritizes early layers for hallucination detection, mid-level layers for OCR, and more diverse mixtures for open-ended reasoning. This pattern indicates that dynamic fusion provides both adaptivity and interpretability.
Although TGIF substantially improves hallucination robustness and fine-grained perception, several challenges remain. First, our fusion mechanism operates over fixed CLIP features; while dynamic routing improves layer selection, the model remains constrained by the representational biases and resolution limits of the frozen encoder. Future work may explore pairing TGIF with higher-resolution or task-specific encoders, or integrating lightweight vision-side adapters to further enhance fine-grained cues. Second, TGIF routes which layers to use but not what spatial regions within each layer to emphasize. Combining inter-layer fusion with adaptive spatial attention or region-level routing may further strengthen grounding in cluttered or text-heavy scenes.
Overall, TGIF demonstrates that hierarchy-aware visual fusion is a promising and scalable path toward more grounded, trustworthy, and task-aware multimodal LLMs, while also opening opportunities for deeper exploration of dynamic visual representations.
[…]
Supplementary Material
6 Implementation Details
Hardware & Precision. All experiments are conducted on a single node equipped with $8\times$ NVIDIA H100 (80GB) GPUs. We utilize DeepSpeed ZeRO (Stage 2 for pretraining, Stage 3 for instruction tuning) to optimize memory efficiency. Unless otherwise specified, we employ Vicuna-7B-v1.5 as the language model and CLIP ViT-L/14@336px as the frozen vision encoder. Training is performed with bf16 precision and tf32 matrix multiplication acceleration.
Stage 1: Feature Alignment. Following the LLaVA-1.5 protocol, we train only the multimodal projector (including the proposed TGIF router) for 1 epoch on the LLaVA pretraining dataset (558K image–text pairs). We utilize a learning rate of $1\times 10^{-3}$ with a cosine decay schedule and a warmup ratio of 0.03. The global batch size is set to 32 (gradient accumulation steps =1=1), with a maximum sequence length of 2048. Weight decay is set to 0.
Stage 2: Instruction Tuning. In the second stage, we fine-tune the projector and the LLM on the LLaVA v1.5 mixture (665K samples) for 1 epoch. We reduce the learning rate to $2\times 10^{-5}$ while maintaining the cosine schedule, warmup ratio of 0.03, and weight decay of 0. The per-GPU batch size is adjusted to 16.
Throughout both stages, the vision encoder remains entirely frozen. We enable gradient checkpointing to conserve memory. All specific hyperparameters and script templates are provided in the attached code for reproducibility.
7 Benchmark Descriptions
7.1 Hallucination Evaluation
POPE [Li-hallucination-2023].
POPE assesses object hallucination via a binary verification task. For a given image, the model must answer Yes/No questions (e.g., “Is there a <object> in the image?”). Negative samples are generated using three strategies: (1) Random sampling, (2) Popular COCO categories, and (3) Adversarial co-occurring objects. This setup isolates visual grounding capabilities from captioning priors. We report Accuracy, Precision, Recall, F1-score, and the ”Yes” response ratio.
HallusionBench [hallusionbench]
This benchmark evaluates visual factual grounding by presenting questions with fabricated or contradictory premises. Unlike standard VQA, HallusionBench penalizes models for failing to reject incorrect visual claims regarding object existence, attributes, and spatial relations. Following the official protocol, we employ a GPT-4 assisted evaluation metric. We report aAcc (All Accuracy) as the primary metric, alongside qAcc (Question-Pair Accuracy) and fAcc (Figure Accuracy).
7.2 OCR Evaluation
OCRBench [Liu_2024].
OCRBench is a comprehensive evaluation suite consisting of 1,000 manually verified QA pairs aggregated from 29 diverse datasets. It assesses five core capabilities: Text Recognition, Scene Text VQA, Document VQA, Key Information Extraction (KIE), and Handwritten Math Expression Recognition (HMER). Evaluation is performed via exact string matching.
8 Additional Quantitative Results
8.1 POPE Breakdown
In Table 5, we provide a detailed breakdown of performance across the Random, Popular, and Adversarial splits of the POPE benchmark. Our method (TGIF) consistently improves Precision and Accuracy across all subsets compared to the LLaVA-1.5 baseline, indicating robust resistance to hallucination regardless of the negative sampling strategy.
8.2 HallusionBench Leaderboard
Table 6 presents the full HallusionBench correctness leaderboard. TGIF achieves competitive performance among 7B parameters models, particularly in the aAcc metric, surpassing the LLaVA-1.5 baseline and approaching the performance of proprietary models like Gemini Pro Vision.
Notably, despite utilizing a significantly smaller language backbone (7B parameters), TGIF (49.94%) outperforms several larger open-source baselines, including the 13B-parameter LLaVA-1.5 (46.94%) and the 12.1B-parameter BLIP2-T5 (48.09%) [liu2023improved, Li2023BLIP2BL]. It secures the third rank overall, trailing only the closed-source models GPT-4V and Claude 3 [2023GPT4VisionSC, claude3]. This result shows the efficiency of our text-guided fusion strategy: by dynamically routing to the most relevant visual features, TGIF extracts rich grounding signals from the frozen encoder, effectively mitigating hallucination even against models with nearly double the parameter count.
8.3 OCRBench Performance
We report detailed OCRBench scores in Table 7. TGIF demonstrates improvements in Scene Text VQA ($VQA^{S}$) and Document VQA ($VQA^{D}$), contributing to a higher final score compared to the LLaVA-1.5 baseline. This suggests that layer fusion effectively captures fine-grained textual details often lost in single-layer embeddings.
Table 5: POPE Results by Subset. Comparison of LLaVA-1.5-7B vs. LLaVA-1.5-7B+TGIF across specific hallucination settings. Best performance per subset is bold.
Subset Method F1 Acc Prec Rec Yes %
Random LLaVA-1.5 0.873 0.882 0.975 0.791 41.9
+ TGIF 0.891 0.895 0.960 0.831 44.6
Popular LLaVA-1.5 0.861 0.872 0.944 0.791 41.9
+ TGIF 0.876 0.882 0.926 0.831 44.9
Adversarial LLaVA-1.5 0.842 0.851 0.899 0.791 44.0
+ TGIF 0.856 0.860 0.882 0.831 47.1
Table 6: HallusionBench Correctness Leaderboard. We report Question-pair Accuracy (qAcc), Figure Accuracy (fAcc), and All Accuracy (aAcc). Top-3 models under the GPT-4–assisted evaluation are highlighted in bold.
Method Params Eval Mode qAcc $\uparrow$ fAcc $\uparrow$ aAcc $\uparrow$
GPT-4V [2023GPT4VisionSC] (Oct 2023) - Human 31.42 44.22 67.58
GPT-4 28.79 39.88 65.28
LLaVA-1.5 [liu2023improved] 13B Human 9.45 25.43 47.12
GPT-4 10.55 24.86 46.94
Claude 3 [claude3] - GPT-4 21.76 28.61 56.86
Gemini Pro Vision [google_gemini] - GPT-4 7.69 8.67 36.85
BLIP2-T5 [Li2023BLIP2BL] 12.1B GPT-4 15.16 20.52 48.09
Qwen-VL [Bai2023QwenVLAF] 9.6B GPT-4 5.93 6.65 39.15
Open-Flamingo [alayrac2022flamingo] 9B GPT-4 6.37 11.27 38.44
MiniGPT-5 [Zheng2023MiniGPT5IV] 8.2B GPT-4 10.55 9.83 40.30
MiniGPT-4 [zhu2023minigpt] 8.2B GPT-4 8.79 10.12 35.78
InstructBLIP [dai2023instructblip] 8.2B GPT-4 9.45 10.11 45.26
BLIP-2 [Li2023BLIP2BL] 8.2B GPT-4 5.05 12.43 40.48
mPLUG-Owl v2 [ye2023mplugowl2] 8.2B GPT-4 13.85 19.94 47.30
mPLUG-Owl v1 [ye2023mplug] 7.2B GPT-4 9.45 10.40 43.93
LRV-Instruction [liu2023aligning] 7.2B GPT-4 8.79 13.01 42.78
TGIF (Ours) 7B GPT-4 17.36 23.70 49.94
GIT [Wang2022GITAG] 0.8B GPT-4 5.27 6.36 34.37
Random Chance - GPT-4 15.60 18.21 45.96
Table 7: Detailed Results on OCRBench. Breakdown of sub-tasks: Text Recognition (Recog.), Scene Text VQA ($VQA^{S}$), Document VQA ($VQA^{D}$), Key Information Extraction (KIE), and Handwritten Math (HMER). Best results are marked in bold.
Method Recog. $VQA^{S}$ $VQA^{D}$ KIE HMER Total
Gemini Pro [google_gemini] 215 174 128 134 8 659
GPT-4V [2023GPT4VisionSC] 167 163 146 160 9 645
mPLUG-Owl2 [ye2023mplugowl2] 153 153 41 19 0 366
LLaVAR [Zhang2023LLaVAREV] 186 122 25 13 0 346
LLaVA-1.5-13B [liu2023improved] 176 129 19 7 0 331
LLaVA-1.5-7B [liu2023improved] 160 117 15 5 0 297
TGIF (Ours) 162 121 24 6 0 313
mPLUG-Owl [mplug-owl] 172 104 18 3 0 297
InstructBLIP [instructblip] 168 93 14 1 0 276
BLIP-2 [Li2023BLIP2BL] 154 71 10 0 0 235
MiniGPT-4 v2[minigpt-4] 124 29 4 0 0 157